# Atmospheric Mass Loss from TOI-1259 A b, a Gas Giant Planet With a White Dwarf Companion

### v1

@@ 4.2 TOI-1259 A b is not significantly affected by photoevaporation
Using our retrieved mass loss rates, we estimate the planet’s cumulative maximum mass loss by first calculating a mass loss efficiency (the efficiency with which the atmosphere converts incident radiation into heat). Following the procedure outlined in Vissapragada et al. (2022a), and adopting our maximum mass loss rate ( g s-1, measured with Palomar/WIRC) we obtain a heating efficiency of 0.063. Assuming this mass loss efficiency is constant, we calculate the maximum fraction of the planet mass that has been lost to photoevaporation following the approach detailed in Vissapragada et al. (2022a). In order to calculate, the maximum fraction of the planet’s mass lost, we adopt an integrated XUV flux of erg from the initial fast rotator track of Johnstone et al. (2021) as an upper bound on the XUV flux. We find that TOI-1259 A b has lost 0.51 of its initial mass to photoevaporation. We note that even for an order of magnitude larger mass loss rate, the planet still loses 5% of its initial mass to photoevaporation. These calculations assume that the planet either formed in situ, or migrated very early (i.e., before the dissipation of the gas disk). The unique architecture of this system instead indicates that this planet could have migrated in relatively late via the Kozai mechanism (see Sec. 4.3). If so, its integrated total mass lost to date should be even lower.
This calculation tells us about the planet’s past mass loss history. However, we can also look forward in time and calculate whether or not TOI-1259 A b will survive over the remaining main sequence lifetime of it’s host star, assuming a fixed mass loss rate of g s-1 (we note that this is likely an overestimate as the stellar activity of the star should decrease with age, reducing ). We find that it would take 1,670 Gyr for photoevaporation to strip the planet of its envelope, and conclude that the planet should remain stable against photoevaporation over the star’s remaining main sequence lifetime. We also note that our estimate is comparable to the predicted atmospheric lifetimes of other gas-giant planets undergoing photoevaporation near the upper edge of the Neptune desert (Vissapragada et al., 2022b).
@@ 5 Conclusions
In this work, we present the first detection of an escaping atmosphere on TOI-1259 A b, a Saturn-mass planet orbiting a K dwarf with a white dwarf companion. Using Palomar/WIRC we measure an excess metastable helium absorption signal in the transit light curve of . Using Keck/NIRSPEC we measure a blue-shifted excess absorption of , which we integrate with the Palomar/WIRC helium filter to get an excess absorption of 0.26%, which is consistent with the Palomar/WIRC signal within 1.5. We convert the Palomar/WIRC and Keck/NIRSPEC excess absorption measurements into constraints on the mass loss rate of the outflow and find that the planet is losing mass at a rate of g s-1 with our Palomar/WIRC measurement, and g s-1 with our Keck/NIRSPEC measurement. Using our maximum mass loss rate, we derive a corresponding predicted atmospheric lifetime () of which is substantially longer than the lifetime of an average main sequence star. We also calculate the fraction of the initial planet mass lost to photoevaporation, assuming a constant mass loss efficiency, and find that the planet lost only 0.51% of its initial mass. Therefore the planet does not appear to be significantly sculpted by photoevaporation. This is consistent with results for other planets along the upper edge of the Neptune desert, which are also stable against photoevaporation and display similar mass loss rates and heating efficiencies (Vissapragada et al., 2022b; Dos Santos, 2022).
@@
If TOI-1259 A b formed in a more distant formation location, evidence for this should also be recorded within TOI-1259 A b’s atmosphere. Because HEM can only operate after the gas disk has dissipated, we can expect that TOI-1259 A b’s atmosphere should have been minimally affected by migration-driven accretion and should therefore preserve a pristine record of its formation location. As discussed in Öberg et al. (2011), Madhusudhan (2012), and Chachan et al. (2023) the abundance of refractory elements (e.g. carbon-to-oxygen ratio (C/O)) in the planet’s atmosphere can be used as a tracer of formation location. If TOI-1259 A b formed outside of the water snowline, its atmosphere would have been enriched in oxygen and enhanced by the accretion of small water-rich solids, therefore resulting in a substellar C/O ratio. If TOI-1259 A b instead formed interior to the water ice line, it would have accreted more carbon-rich solids and other refractory elements such as sulfur, which would be reflected in an enhanced refractory element abundance (Turrini et al., 2021; Lothringer et al., 2021; Pacetti et al., 2022; Crossfield, 2023; Tsai et al., 2023). TOI-1259 A b has an equilibrium temperature of 917 K, and we therefore expect that its atmosphere may contain detectable quantities of H2O, CO, CO2, SO2 and CH4. If we can measure the abundances of most or all of these species in its atmosphere using transmission spectroscopy, we can potentially determine where in the disk TOI-1259 A b formed. TOI-1259 A b has a Transmission Spectroscopy Metric (TSM; Kempton et al., 2018) of 184, one of the highest TSMs for a Jupiter-sized planet cooler than 1000 K, rendering TOI-1259 A b a compelling target for future atmospheric characterization studies.
@@ acknowledgments
We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. The specific observations analyzed can be accessed via https:/doi.org/10.17909/t9-nmc8-f686 (catalog https:/doi.org/10.17909//t9-nmc8-f686). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE‐1745301. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.

### v2

@@ 4.2 TOI-1259 A b is not significantly affected by photoevaporation
Using our retrieved mass loss rates, we estimate the planet’s cumulative maximum mass loss by first calculating a mass loss efficiency (the efficiency with which the atmosphere converts incident radiation into heat). Following the procedure outlined in Vissapragada et al. (2022a), and adopting our maximum mass loss rate using HD 40307 ( g s-1) and Eridani ( g s-1) we obtain a heating efficiency of 0.075 (HD 40307) and 0.054 ( Eridani). Assuming this mass loss efficiency is constant, we calculate the maximum fraction of the planet mass that has been lost to photoevaporation following the approach detailed in Vissapragada et al. (2022a). In order to calculate, the maximum fraction of the planet’s mass lost, we adopt an integrated XUV flux of erg from the initial fast rotator track of Johnstone et al. (2021) as an upper bound on the XUV flux. Repeating our calculations for both maximum mass loss rates and heating efficiencies, we find that TOI-1259 A b has lost at most 0.61 of its initial mass to photoevaporation. These calculations assume that the planet either formed in situ, or migrated very early (i.e., before the dissipation of the gas disk). The unique architecture of this system instead indicates that this planet could have migrated in relatively late via the Kozai mechanism (see Sec. 4.3). If so, its integrated total mass lost to date should be even lower.
This calculation tells us about the planet’s past mass loss history. However, we can also look forward in time and calculate whether or not TOI-1259 A b will survive over the remaining main sequence lifetime of it’s host star, assuming a fixed mass loss rate of g s-1 (HD 40307) and g s-1 ( Eridani) (we note that this is likely an overestimate as the stellar activity of the star should decrease with age, reducing ). We find that it would take between 594-1,060 Gyr for photoevaporation to strip the planet of its envelope, and conclude that the planet should remain stable against photoevaporation over the star’s remaining main sequence lifetime. Our estimate is comparable to the predicted atmospheric lifetimes of other gas-giant planets undergoing photoevaporation near the upper edge of the Neptune desert (Vissapragada et al., 2022b).
Recent studies (e.g. Lampón et al., 2023) have proposed that some escaping atmospheres can have super-solar hydrogen to helium ratios (e.g., 99/1 H/He), potentially resulting in mass loss rates that are an order of magnitude larger for the same helium absorption signal. We evaluate the effect of a 99/1 H/He atmospheric composition on our estimated mass loss rate for TOI-1259 A b by running a Parker wind model with this alternative composition. We find a g s-1, which is an order of magnitude larger than our mass loss rate for a 90/10 H/He atmosphere, consistent with the predictions of Lampón et al. (2023). Even with this higher mass loss rate, we find that the planet lost of its initial mass and it would take approximately 58 Gyr to completely strip the planet’s atmosphere. We conclude that even if the outflow is significantly enhanced in hydrogen relative to helium, the planet should still retain its atmosphere over the main sequence lifetime of the host star.
@@ 5 Conclusions
In this work, we present the first detection of an escaping atmosphere on TOI-1259 A b, a Saturn-mass planet orbiting a K dwarf with a white dwarf companion. Using Palomar/WIRC we measure an excess metastable helium absorption signal in the transit light curve of . Using Keck/NIRSPEC we measure a blue-shifted peak excess absorption of between first and fourth contact and between second and third contact. Integrating our Keck/NIRSPEC signals with the Palomar/WIRC bandpass, we find that our second and third contact signal, which is more reflective of the behavior at mid-transit, is consistent with the Palomar/WIRC mid-transit excess absorption at the 1 level. We convert the Palomar/WIRC and Keck/NIRSPEC excess absorption measurements into constraints on the mass loss rate of the outflow and find that the planet is losing mass at a rate of g s-1 with our Palomar/WIRC measurement, and g s-1 () and g s-1 () with our Keck/NIRSPEC measurements. However, a single epoch of HIRES observations reveals a potentially active host star, so we repeat our mass loss calculations with a more active proxy stellar spectrum and find that at most, the planet is losing mass at a rate of g s-1. Using our maximum mass loss rate, we derive a corresponding predicted atmospheric lifetime () of which is substantially longer than the lifetime of an average main sequence star. We also calculate the fraction of the initial planet mass lost to photoevaporation, assuming a constant mass loss efficiency, and find that the planet lost at most, only 0.61% of its initial mass. Therefore the planet does not appear to be significantly sculpted by photoevaporation. This is consistent with results for other planets along the upper edge of the Neptune desert, which are also stable against photoevaporation and display similar mass loss rates and heating efficiencies (Vissapragada et al., 2022b; Dos Santos, 2022).
@@
If TOI-1259 A b formed in a more distant formation location, evidence for this should also be recorded within TOI-1259 A b’s atmosphere. Because HEM can only operate after the gas disk has dissipated, we can expect that TOI-1259 A b’s atmosphere should have been minimally affected by migration-driven accretion and should therefore preserve a pristine record of its formation location. As discussed in Öberg et al. (2011), Madhusudhan (2012), and Chachan et al. (2023) the abundance of refractory elements (e.g. the carbon-to-oxygen ratio C/O) in the planet’s atmosphere can be used as a tracer of formation location. If TOI-1259 A b formed outside of the water snowline, its atmosphere would have been enriched in oxygen and enhanced by the accretion of small water-rich solids, therefore resulting in a substellar C/O ratio. If TOI-1259 A b instead formed interior to the water ice line, it would have accreted more carbon-rich solids and other refractory elements such as sulfur, which would be reflected in an enhanced refractory element abundance (Turrini et al., 2021; Lothringer et al., 2021; Pacetti et al., 2022; Crossfield, 2023; Tsai et al., 2023). TOI-1259 A b has an equilibrium temperature of 917 K, and we therefore expect that its atmosphere may contain detectable quantities of H2O, CO, CO2, SO2 and CH4. If we can measure the abundances of most or all of these species in its atmosphere using transmission spectroscopy, we can potentially determine where in the disk TOI-1259 A b formed. TOI-1259 A b has a Transmission Spectroscopy Metric (TSM; Kempton et al., 2018) of 184, one of the highest TSMs for a Jupiter-sized planet cooler than 1000 K, rendering TOI-1259 A b a compelling target for future atmospheric characterization studies.
@@ acknowledgments
We acknowledge the use of public TESS data from pipelines at the TESS Science Office and at the TESS Science Processing Operations Center. This paper includes data collected by the TESS mission that are publicly available from the Mikulski Archive for Space Telescopes (MAST). Funding for the TESS mission is provided by NASA’s Science Mission Directorate. The specific observations analyzed can be accessed via https:/doi.org/10.17909/t9-nmc8-f686 (catalog https:/doi.org/10.17909//t9-nmc8-f686) (MAST Team, 2021). STScI is operated by the Association of Universities for Research in Astronomy, Inc., under NASA contract NAS5–26555. Support to MAST for these data is provided by the NASA Office of Space Science via grant NAG5–7584 and by other grants and contracts. This research has made use of the NASA Exoplanet Archive, which is operated by the California Institute of Technology, under contract with the National Aeronautics and Space Administration under the Exoplanet Exploration Program. This material is based upon work supported by the National Science Foundation Graduate Research Fellowship Program under Grant No. DGE‐1745301. Any opinions, findings, and conclusions or recommendations expressed in this material are those of the author(s) and do not necessarily reflect the views of the National Science Foundation.